Allison:

Setting up the proportion that indicates the percent of change, without solving~ Answer checking.

Allison:

From 59 to 99 40/59 = % change/100

Hero:

Looks right :)

Allison:

:D

Allison:

From 59 to 45 -14/59 = % change/100

Hero:

But remember the formula I gave you in which I said we only include the positive difference between the two numbers?

Allison:

Oh yes

Allison:

Whoops o.o

Allison:

14/59 :D

Allison:

From 11 to 5 6/11 = % change/100

Hero:

Correct. However, it seems you're getting the hang of it now.

Allison:

Yes thank you for the reminder for the negative and positive formula.

Allison:

Is it the same with decimals..?

Allison:

@Hero